OpenStudy (anonymous):

Choose the correct ordered pair. 3x - y = 3 5x + 3y = -23

OpenStudy (anonymous):

Options?

OpenStudy (anonymous):

Its a choice right?

OpenStudy (callisto):

3x - y = 3 => y = 3x-3 -(1) 5x + 3y = -23 - (2) Sub (1) into (2) 5x + 3(3x-3 ) = -23 Can you solve x from here?

OpenStudy (anonymous):

A. (1, 6) B. (1, -6) C. (-1, -6) D. (-1, 6)

OpenStudy (callisto):

After finding x, put x =? into y = 3x-3 , you'll get the y value

OpenStudy (anonymous):

Multiply the first equation by 3 and add it to the second equation.Now solve for x and substitute this value in any one of the two equations and get y.

OpenStudy (callisto):

@daijagattison Can you solve 5x + 3(3x-3 ) = -23 ?

OpenStudy (anonymous):

(-1, 6)

OpenStudy (callisto):

Nope... x is correct, y is incorrect

OpenStudy (anonymous):

(-1, -6)

OpenStudy (callisto):

Yes!